5. Use the Laplace transform to solve the following initial value problem: y - 6y +9y = 3e-21, y(0) = 1, y(0) = -1.
Answer #1

y eyt ay = ze at > d Cy) - 6 d (y] +9.dcy)=3 d Ce- 28 dCy) – 2yco)- ylio) -6 [8d04) – you] ta dicy) = 3 - eq (1) 1 we also

Alat €26510 1e (376549)(5+2) S s _3_ + S (5-3)&25+23 (3- = 3+32+2 8 = 48-19 (S-332 estaj a $2.55 - 11 : (5-3)2C5+23 5-3 (9-3)

- (837 € C20 - Casa St 5-3 5 (5-3) nga dalam 1()-921654), 2(4) :., ten2014 G Cans) Note z Juleat) di ceeat) - saja